# Enhanced simultaneous quantum-classical communications under composable security

### v1

@@ II.1 General formulation of SQCC protocols in the prepare-and-measure basis
    Alice begins by randomly drawing some value $\tilde{\alpha}$ from a zero-mean bivariate Gaussian distribution of variance $V$, and prepares the coherent state |$\ket{\tilde{\alpha}}$. This represents the quantum symbol used for key generation.
@@ II.2 Security analysis
$\bm{V}$ $=\begin{pmatrix}V\mathbb{I}&\sqrt{V^{2}-1}\sigma_{Z}\\ \sqrt{V^{2}-1}\sigma_{Z}&V\mathbb{I}\end{pmatrix},$ (1)
@@
After the channel, Alice and Bob share a mixed two-mode entangled Gaussian state fully described by a mean vector $\bm{\mu}$ and covariance matrix $\bm{V}$ [22]:
@@
$\bm{V}$ $=\begin{pmatrix}V\mathbb{I}&\sqrt{T(V^{2}-1)}\sigma_{Z}\\ \sqrt{T(V^{2}-1)}\sigma_{Z}&[T(V+\varepsilon-1)+1]\mathbb{I}\end{pmatrix}$ (4)
@@
Alice and Bob now perform heterodyne detection on their respective modes and obtain complex measurement outcomes $\tilde{\alpha}\equiv\alpha_{x}+i\alpha_{y}$ and $\tilde{\beta}\equiv\beta_{x}+i\beta_{y}$ respectively. By construction, their measurement results are distributed according to the Husimi Q-function of the state described by the mean vector $\bm{\mu}$ and covariance matrix $\bm{V}$ given in Eqs. (34). Their joint measurement ensemble therefore has the probability density function (Appendix A)
$Q(\tilde{\alpha},\tilde{\beta})=Q^{\bm{\mu}}_{\bm{V}}(\alpha_{x},\alpha_{y},\beta_{x},\beta_{y}).$ (6)
@@
$Q_{in}(\bm{\alpha})$ $=\frac{Q^{d_{1}}+Q^{d_{2}}+Q^{d_{3}}+Q^{d_{4}}}{4}.$ (7)
We might also justify this on a shot-by-shot basis by recalling that Bob does not know which of the four symbols he is going to receive, and so the quantum state he receives is a probabilistic mix of the four encodings. (Here we use the shorthand $Q^{d_{k}}$ to indicate the Q-function $Q^{\bm{\mu}}_{\bm{V}}$ of the state with covariance $\bm{V}$ and mean $\bm{\mu}$ associated with the classical displacement $\tilde{d}=d_{k}$ as per Eqs. [34]).
@@ III.1 Monte Carlo simulation of signal postprocessing
[…]
@@
$\bm{V}$ $=\begin{pmatrix}V\mathbb{I}&\sqrt{T(V^{2}-1)}\sigma_{Z}\\ \sqrt{T(V^{2}-1)}\sigma_{Z}&\left[T(V+\varepsilon^{\prime}-1)+1\right]\mathbb{I}\end{pmatrix},$ (8)
@@ III.2 Analytic form of the covariance matrix
$\bm{\mu}$ $=\left(0,0,\frac{\alpha}{\sqrt{2}},\frac{\alpha}{\sqrt{2}}\right)^{T},$ (10)
$\bm{V}$ $=\begin{pmatrix}a\mathbb{I}&c\sigma_{Z}\\ c\sigma_{Z}&b\mathbb{I}\end{pmatrix}$ (11)
@@
$\bm{\mu}_{d}$ $=\left(0,0,\frac{2\alpha e_{C}}{\sqrt{2}},\frac{2\alpha e_{C}}{\sqrt{2}}\right)^{T},$ (13)
$\bm{V}_{d}$ $=\begin{pmatrix}a_{d}\mathbb{I}&c_{d}\sigma_{Z}\\ c_{d}\sigma_{Z}&b_{d}\mathbb{I}\end{pmatrix}$ (14)
@@ III.3 Renormalisation of the postprocessed state
$\bm{\mu}_{d}^{\prime}$ $=\left(0,0,\frac{1}{\sqrt{\Delta_{V}}}\frac{2\alpha e_{C}}{\sqrt{2}},\frac{1}{\sqrt{\Delta_{V}}}\frac{2\alpha e_{C}}{\sqrt{2}}\right)^{T},$ (21)
$\bm{V}_{d}{{}^{\prime}}$ $=\begin{pmatrix}a_{d}\mathbb{I}&\frac{c_{d}}{\sqrt{\Delta_{V}}}\sigma_{Z}\\ \frac{c_{d}}{\sqrt{\Delta_{V}}}\sigma_{Z}&\left[\frac{b_{d}}{\Delta_{V}}+\frac{1}{\Delta_{V}}-1\right]\mathbb{I}\end{pmatrix}.$ (22)
@@
$\bm{V}_{d}{{}^{\prime}}$ $=\begin{pmatrix}a\mathbb{I}&c\sigma_{Z}\\ c\sigma_{Z}&\left[\frac{b_{d}+1}{(1+\delta^{2})}-1\right]\mathbb{I}\end{pmatrix}$ (24)
@@
$\bm{V}_{d}{{}^{\prime}}$ $=\begin{pmatrix}a\mathbb{I}&c\sqrt{\frac{b+1}{b_{d}+1}}\sigma_{Z}\\ c\sqrt{\frac{b+1}{b_{d}+1}}\sigma_{Z}&b\mathbb{I}\end{pmatrix}$ (28)
@@ III.5 Finite-key regime
Further technical details on the parameter estimation process are presented in Appendix D. The maximum secret-key bits shared per pulse $K^{\mathcal{F}}\leavevmode\nobreak\ =\leavevmode\nobreak\ \ell/N$ of such a protocol is therefore bounded below by
@@ Appendix A Husimi Q-function formulation
$Q(\bm{\alpha})$ $=\frac{1}{\pi^{n}}\bra{\bm{\alpha}}\hat{\rho}\ket{\bm{\alpha}},$ (43)
where $\bm{\alpha}\in\mathbb{C}^{n}$ and |$\ket{\bm{\alpha}}\equiv\bigotimes_{i=0}^{n}\ket{\alpha_{i}}$. By construction, the Q-function is the normalised and positive probability density function [28] associated with projection onto a specific coherent state |$\ket{\bm{\alpha}}$, which also implies it is the probability density function associated with Alice and Bob’s measurement results for their protocol when using heterodyne measurement [20].
Since our protocol deals commonly with Gaussian states, it is useful to derive a general formula for the Q-function of a Gaussian state given an arbitrary mean vector and covariance matrix. Using shot-noise units (=2Planck-constant-over-$\hbar=2$) [22], the Wigner function of a general $n$-mode Gaussian state is given by [20]
$W^{\bm{\mu}}_{\bm{V}}(\bm{\xi})$ $=\frac{e^{-\frac{1}{2}(\bm{\xi}-\bm{\mu})^{T}\bm{V}^{-1}(\bm{\xi}-\bm{\mu})}}{(2\pi)^{n}\sqrt{\det(\bm{V})}}$ (44)
for ${\bm{\mu}}=(\langle\hat{q}_{A}\rangle,\langle\hat{p}_{A}\rangle,...)^{T}$ the mean of the quantum state, $\bm{\xi}\in\mathbb{R}^{2n}$ and $\bm{V}\in\text{Sp}(2n,\mathbb{R})$ the symplectic covariance matrix. We obtain the Husimi Q-function from the Wigner function via Weierstrass transform:
$Q(\bm{\alpha})$ $=\bigg{(}\frac{2}{\pi}\bigg{)}^{n}\int\text{d}\bm{\xi}\ W(\bm{\xi})e^{-\frac{1}{2}(2\bm{\alpha}-\bm{\xi})^{T}(2\bm{\alpha}-\bm{\xi})}.$ (45)
It is not difficult to show then that the generalised Q-function of a Gaussian state with mean vector $\bm{\mu}$ and covariance matrix $\bm{V}$ is given by
$Q^{\bm{\mu}}_{\bm{V}}(\bm{\alpha})=\frac{\sqrt{\det(\bm{M})}}{\pi^{N}}e^{-(\bm{\alpha}-\frac{\bm{\mu}}{2})^{T}\bm{M}(\bm{\alpha}-\frac{\bm{\mu}}{2})},$ (46)
for $\bm{M}=2(\bm{V}+\mathbb{I})^{-1}$, when $V$ has the general form given by Eq. (4). (This is consistent with the zero-mean expression given in [29].) Finally, the statistical moments of the quadrature operators in the quantum space are functions of the statistical moments of the Q-function in the phase space:
@@
A useful property of the Husimi Q-function is that for a Gaussian quantum state of mean $\bm{\mu}$ and covariance $\bm{V}$, the Q-function is itself Gaussian in $\bm{\alpha}$ with mean $\bm{\mu}/2$ and covariance $(\bm{V}+\mathbb{I})/4$.
@@ Appendix B Derivation of the covariance matrix
$\bm{\mu}$ $=\left(0,0,\frac{\sqrt{T}d}{\sqrt{2}},\frac{\sqrt{T}d}{\sqrt{2}}\right)^{T},$ (50)
$\bm{V}$ $=\begin{pmatrix}V\mathbb{I}&\sqrt{T(V^{2}-1)}\sigma_{Z}\\ \sqrt{T(V^{2}-1)}\sigma_{Z}&\left[T(V+\varepsilon-1)+1\right]\mathbb{I}\end{pmatrix}$ (51)
@@ B.1 Measurement statistics after postprocessing
$Q_{in}=\begin{cases}Q^{d_{1}}(\tilde{\alpha},\tilde{\beta})&\beta_{x}\geq 0,\beta_{y}\geq 0\ (d_{1},00)\\ Q^{d_{1}}(\tilde{\alpha},\tilde{\beta})&\beta_{x}<0,\beta_{y}>0\ (d_{2},01)\\ Q^{d_{1}}(\tilde{\alpha},\tilde{\beta})&\beta_{x}\leq 0,\beta_{y}\leq 0\ (d_{3},11)\\ Q^{d_{1}}(\tilde{\alpha},\tilde{\beta})&\beta_{x}>0,\beta_{y}<0\ (d_{4},10).\\ \end{cases}$ (55)
He then virtually performs the re-displacement operation $\hat{D}(-\sqrt{T}\tilde{d})$ to each shot according to the measured classical symbol to return the distribution to a zero-mean one, approximately reproducing the original quantum signal. This is equivalent to the transform $\tilde{\beta}\Longrightarrow\tilde{\beta}^{\prime}=\tilde{\beta}-\sqrt{T}\tilde{d}$, and so the Q-function $Q_{d}$ describing the distribution of Alice and Bob’s re-displaced quantum measurement results becomes
$Q_{d}$ =$=\begin{cases}Q^{d_{1}}(\tilde{\alpha},\tilde{\beta}^{\prime}+\sqrt{T}d_{1})&\beta_{x}^{\prime}+\frac{\sqrt{T}d}{2\sqrt{2}}\geq 0,\beta_{y}^{\prime}+\frac{\sqrt{T}d}{2\sqrt{2}}\geq 0\\ Q^{d_{1}}(\tilde{\alpha},\tilde{\beta}^{\prime}+\sqrt{T}d_{2})&\beta_{x}^{\prime}-\frac{\sqrt{T}d}{2\sqrt{2}}<0,\beta_{y}^{\prime}+\frac{\sqrt{T}d}{2\sqrt{2}}>0\\ Q^{d_{1}}(\tilde{\alpha},\tilde{\beta}^{\prime}+\sqrt{T}d_{3})&\beta_{x}^{\prime}-\frac{\sqrt{T}d}{2\sqrt{2}}\leq 0,\beta_{y}^{\prime}-\frac{\sqrt{T}d}{2\sqrt{2}}\leq 0\\ Q^{d_{1}}(\tilde{\alpha},\tilde{\beta}^{\prime}+\sqrt{T}d_{4})&\beta_{x}^{\prime}+\frac{\sqrt{T}d}{2\sqrt{2}}>0,\beta_{y}^{\prime}-\frac{\sqrt{T}d}{2\sqrt{2}}<0\\ \end{cases}$ (56)
=$=\begin{cases}Q^{0}(\tilde{\alpha},\tilde{\beta}^{\prime})&\beta_{x}^{\prime}\geq-\frac{\sqrt{T}d}{2\sqrt{2}},\ \beta_{y}^{\prime}\geq-\frac{\sqrt{T}d}{2\sqrt{2}}\\ Q^{d_{1}-d_{2}}(\tilde{\alpha},\tilde{\beta}^{\prime})&\beta_{x}^{\prime}<\frac{\sqrt{T}d}{2\sqrt{2}},\ \ \ \beta_{y}^{\prime}>-\frac{\sqrt{T}d}{2\sqrt{2}}\\ Q^{d_{1}-d_{3}}(\tilde{\alpha},\tilde{\beta}^{\prime})&\beta_{x}^{\prime}\leq\frac{\sqrt{T}d}{2\sqrt{2}},\ \ \ \beta_{y}^{\prime}\leq\frac{\sqrt{T}d}{2\sqrt{2}}\\ Q^{d_{1}-d_{4}}(\tilde{\alpha},\tilde{\beta}^{\prime})&\beta_{x}^{\prime}>-\frac{\sqrt{T}d}{2\sqrt{2}},\ \beta_{y}^{\prime}<\frac{\sqrt{T}d}{2\sqrt{2}}\\ \end{cases}$ (57)
$\equiv\begin{cases}Q^{0}&\beta_{x}\geq-\frac{\sqrt{T}d}{2\sqrt{2}},\ \beta_{y}\geq-\frac{\sqrt{T}d}{2\sqrt{2}}\\ Q^{2d/\sqrt{2}}&\beta_{x}<\frac{\sqrt{T}d}{2\sqrt{2}},\ \ \ \beta_{y}>-\frac{\sqrt{T}d}{2\sqrt{2}}\\ Q^{2d/\sqrt{2}+2id/\sqrt{2}}&\beta_{x}\leq\frac{\sqrt{T}d}{2\sqrt{2}},\ \ \ \beta_{y}\leq\frac{\sqrt{T}d}{2\sqrt{2}}\\ Q^{2id/\sqrt{2}}&\beta_{x}>-\frac{\sqrt{T}d}{2\sqrt{2}},\ \beta_{y}<\frac{\sqrt{T}d}{2\sqrt{2}}.\\ \end{cases}$ (58)
@@
$\bm{\mu}_{d}$ $=\left(0,0,\frac{2\sqrt{T}de_{C}}{\sqrt{2}},\frac{2\sqrt{T}de_{C}}{\sqrt{2}}\right)^{T},$ (59)
$\bm{V}_{d}$ $=\begin{pmatrix}a_{d}\mathbb{I}&c_{d}\sigma_{Z}\\ c_{d}\sigma_{Z}&b_{d}\mathbb{I}\end{pmatrix}$ (60)
@@
In addition, the postprocessed state $Q_{d}$ remains normalised, i.e. $||Q_{d}||=1$. This is to be expected, as we are not removing any data points or altering the probability ‘volume’ using any methods other than slicing and translation, which leaves the total probability the same as the initial (normalised) state.

### v2

@@ II.1 General formulation of SQCC protocols in the prepare-and-measure basis
[…]
@@ II.2 Security analysis
$\boldsymbol{V}$ $=\begin{pmatrix}V\mathbb{I}&\sqrt{V^{2}-1}\sigma_{Z}\\ \sqrt{V^{2}-1}\sigma_{Z}&V\mathbb{I}\end{pmatrix}$, (1)
@@
After the channel, Alice and Bob share a mixed two-mode entangled Gaussian state fully described by a mean vector $\boldsymbol{\mu}$ and covariance matrix $\boldsymbol{V}$ [22]:
@@
$\boldsymbol{V}$ $=\begin{pmatrix}V\mathbb{I}&\sqrt{T(V^{2}-1)}\sigma_{Z}\\ \sqrt{T(V^{2}-1)}\sigma_{Z}&[T(V+\varepsilon-1)+1]\mathbb{I}\end{pmatrix}$ (4)
@@
Alice and Bob now perform heterodyne detection on their respective modes and obtain complex measurement outcomes $\tilde{\alpha}\equiv\alpha_{x}+i\alpha_{y}$ and $\tilde{\beta}\equiv\beta_{x}+i\beta_{y}$ respectively. By construction, their measurement results are distributed according to the Husimi Q-function of the state described by the mean vector $\boldsymbol{\mu}$ and covariance matrix $\boldsymbol{V}$ given in Eqs. (3, 4). Their joint measurement ensemble therefore has the probability density function (Appendix A)
$Q(\tilde{\alpha},\tilde{\beta})=Q^{\boldsymbol{\mu}}_{\boldsymbol{V}}(\alpha_{x},\alpha_{y},\beta_{x},\beta_{y}).$ (6)
@@
$Q_{in}(\boldsymbol{\alpha})$ $=\frac{Q^{d_{1}}+Q^{d_{2}}+Q^{d_{3}}+Q^{d_{4}}}{4}.$ (7)
We might also justify this on a shot-by-shot basis by recalling that Bob does not know which of the four symbols he is going to receive, and so the quantum state he receives is a probabilistic mix of the four encodings. (Here we use the shorthand $Q^{d_{k}}$ to indicate the Q-function $Q^{\boldsymbol{\mu}}_{\boldsymbol{V}}$ of the state with covariance $\boldsymbol{V}$ and mean $\boldsymbol{\mu}$ associated with the classical displacement $\tilde{d}=d_{k}$ as per Eqs. [3, 4]).
@@ III.1 Monte Carlo simulation of signal postprocessing
To test our model, we therefore generate a random sample of bivariate complex Gaussian data of size $N\penalty 10000\ =\penalty 10000\ 10^{8}$ according to the mean vector and covariance matrix given in Eqs. (3) and (4) respectively, to simulate the correlated heterodyne measurement results obtained by Alice and Bob for the noisy channel $\mathcal{E}(T,\epsilon)$ (Fig. 2[a]). We postprocess the data blindly according to the threshold discrimination and re-displacement operation described in Section II.1 to generate Bob’s emulated zero-mean distribution (Figure 2[b]). From this emulated distribution, we directly calculate the sample bit-error rate $e_{C}$ and covariance matrix elements and compare them with the analytic predictions (Figure 2[c]) made in the following section.
@@
$\boldsymbol{V}$ $=\begin{pmatrix}V\mathbb{I}&\sqrt{T(V^{2}-1)}\sigma_{Z}\\ \sqrt{T(V^{2}-1)}\sigma_{Z}&\left[T(V+\varepsilon^{\prime}-1)+1\right]\mathbb{I}\end{pmatrix}$, (8)
@@ III.2 Analytic form of the covariance matrix
$\boldsymbol{\mu}$ $=\left(0,0,\frac{\alpha}{\sqrt{2}},\frac{\alpha}{\sqrt{2}}\right)^{T},$ (10)
$\boldsymbol{V}$ $=\begin{pmatrix}a\mathbb{I}&c\sigma_{Z}\\ c\sigma_{Z}&b\mathbb{I}\end{pmatrix}$ (11)
@@
$\boldsymbol{\mu}_{d}$ $=\left(0,0,\frac{2\alpha e_{C}}{\sqrt{2}},\frac{2\alpha e_{C}}{\sqrt{2}}\right)^{T},$ (13)
$\boldsymbol{V}_{d}$ $=\begin{pmatrix}a_{d}\mathbb{I}&c_{d}\sigma_{Z}\\ c_{d}\sigma_{Z}&b_{d}\mathbb{I}\end{pmatrix}$ (14)
@@ III.3 Renormalisation of the postprocessed state
$\boldsymbol{\mu}_{d}^{\prime}$ $=\left(0,0,\frac{1}{\sqrt{\Delta_{V}}}\frac{2\alpha e_{C}}{\sqrt{2}},\frac{1}{\sqrt{\Delta_{V}}}\frac{2\alpha e_{C}}{\sqrt{2}}\right)^{T},$ (21)
$\boldsymbol{V}_{d}{{}^{\prime}}$ $=\begin{pmatrix}a_{d}\mathbb{I}&\frac{c_{d}}{\sqrt{\Delta_{V}}}\sigma_{Z}\\ \frac{c_{d}}{\sqrt{\Delta_{V}}}\sigma_{Z}&\left[\frac{b_{d}}{\Delta_{V}}+\frac{1}{\Delta_{V}}-1\right]\mathbb{I}\end{pmatrix}$. (22)
@@
$\boldsymbol{V}_{d}{{}^{\prime}}$ $=\begin{pmatrix}a\mathbb{I}&c\sigma_{Z}\\ c\sigma_{Z}&\left[\frac{b_{d}+1}{(1+\delta^{2})}-1\right]\mathbb{I}\end{pmatrix}$ (24)
@@
$\boldsymbol{V}_{d}{{}^{\prime}}$ $=\begin{pmatrix}a\mathbb{I}&c\sqrt{\frac{b+1}{b_{d}+1}}\sigma_{Z}\\ c\sqrt{\frac{b+1}{b_{d}+1}}\sigma_{Z}&b\mathbb{I}\end{pmatrix}$ (28)
@@ III.5 Finite-key regime
Further technical details on the parameter estimation process are presented in Appendix D. The maximum secret-key bits shared per pulse $K^{\mathcal{F}}\penalty 10000\ =\penalty 10000\ \ell/N$ of such a protocol is therefore bounded below by
@@ Appendix A Husimi Q-function formulation
[…]
Since our protocol deals commonly with Gaussian states, it is useful to derive a general formula for the Q-function of a Gaussian state given an arbitrary mean vector and covariance matrix. Using shot-noise units ($\hbar=2$) [22], the Wigner function of a general $n$-mode Gaussian state is given by [20]
$W^{\boldsymbol{\mu}}_{\boldsymbol{V}}(\boldsymbol{\xi})$ $=\frac{e^{-\frac{1}{2}(\boldsymbol{\xi}-\boldsymbol{\mu})^{T}\boldsymbol{V}^{-1}(\boldsymbol{\xi}-\boldsymbol{\mu})}}{(2\pi)^{n}\sqrt{\det(\boldsymbol{V})}}$ (44)
for ${\boldsymbol{\mu}}=(\langle\hat{q}_{A}\rangle,\langle\hat{p}_{A}\rangle,...)^{T}$ the mean of the quantum state, $\boldsymbol{\xi}\in\mathbb{R}^{2n}$ and $\boldsymbol{V}\in\text{Sp}(2n,\mathbb{R})$ the symplectic covariance matrix. We obtain the Husimi Q-function from the Wigner function via Weierstrass transform:
$Q(\boldsymbol{\alpha})$ $=\bigg(\frac{2}{\pi}\bigg)^{n}\int\text{d}\boldsymbol{\xi}\ W(\boldsymbol{\xi})e^{-\frac{1}{2}(2\boldsymbol{\alpha}-\boldsymbol{\xi})^{T}(2\boldsymbol{\alpha}-\boldsymbol{\xi})}.$ (45)
It is not difficult to show then that the generalised Q-function of a Gaussian state with mean vector $\boldsymbol{\mu}$ and covariance matrix $\boldsymbol{V}$ is given by
$Q^{\boldsymbol{\mu}}_{\boldsymbol{V}}(\boldsymbol{\alpha})=\frac{\sqrt{\det(\boldsymbol{M})}}{\pi^{N}}e^{-(\boldsymbol{\alpha}-\frac{\boldsymbol{\mu}}{2})^{T}\boldsymbol{M}(\boldsymbol{\alpha}-\frac{\boldsymbol{\mu}}{2})},$ (46)
for $\boldsymbol{M}=2(\boldsymbol{V}+\mathbb{I})^{-1}$, when $V$ has the general form given by Eq. (4). (This is consistent with the zero-mean expression given in [29].) Finally, the statistical moments of the quadrature operators in the quantum space are functions of the statistical moments of the Q-function in the phase space:
@@
A useful property of the Husimi Q-function is that for a Gaussian quantum state of mean $\boldsymbol{\mu}$ and covariance $\boldsymbol{V}$, the Q-function is itself Gaussian in $\boldsymbol{\alpha}$ with mean $\boldsymbol{\mu}/2$ and covariance $(\boldsymbol{V}+\mathbb{I})/4$.
@@ Appendix B Derivation of the covariance matrix
$\boldsymbol{\mu}$ $=\left(0,0,\frac{\sqrt{T}d}{\sqrt{2}},\frac{\sqrt{T}d}{\sqrt{2}}\right)^{T},$ (50)
$\boldsymbol{V}$ $=\begin{pmatrix}V\mathbb{I}&\sqrt{T(V^{2}-1)}\sigma_{Z}\\ \sqrt{T(V^{2}-1)}\sigma_{Z}&\left[T(V+\varepsilon-1)+1\right]\mathbb{I}\end{pmatrix}$ (51)
@@ B.1 Measurement statistics after postprocessing
[…]
@@
$\boldsymbol{\mu}_{d}$ $=\left(0,0,\frac{2\sqrt{T}de_{C}}{\sqrt{2}},\frac{2\sqrt{T}de_{C}}{\sqrt{2}}\right)^{T},$ (59)
$\boldsymbol{V}_{d}$ $=\begin{pmatrix}a_{d}\mathbb{I}&c_{d}\sigma_{Z}\\ c_{d}\sigma_{Z}&b_{d}\mathbb{I}\end{pmatrix}$ (60)
@@
In addition, the postprocessed state $Q_{d}$ remains normalised, i.e. $Q_{d}||=1$. This is to be expected, as we are not removing any data points or altering the probability ‘volume’ using any methods other than slicing and translation, which leaves the total probability the same as the initial (normalised) state.